# IP-LLM: Semantic Space Informed Prompt Learning with LLM for Time Series Forecasting

### v1

@@ 4.1 Long-term Forecasting
Setup. For long-term forecasting, we evaluate the effectiveness of IP-LLM over Weather, Electricity, Traffic, and four ETT datasets (i.e., ETTh1, ETTh2, ETTm1, and ETTm2), which have been widely adopted as benchmarking datasets for long-term forecasting tasks. Details of these datasets are shown in Appendix A.3, Table 6. The input time series length is 512 and we evaluate the performance on four different horizons . The evaluation metrics include the mean square error (MSE) and the mean absolute error (MAE).
@@
Results. We compare the forecasting results of IP-LLM to 6 selected baselines in Table 1. Due to the space limitation, the comparisons with other 5 baselines are provided in Appendix B and Table 7. We can observe that LLMs based forecasting methods, i.e., Time-LLM and OFA, generally achieve better performance than other baseline methods. This should be attributed to the prevalent expressibility of LLMs and their associated prompt-tuning and fine-tuning strategies, respectively. Moreover, IP-LLM consistently outperforms Time-LLM and OFA over 7 different datasets. It is particularly worth mentioning that IP-LLM achieves 17% to 25% improvement in terms of average MSE on ETTm1 and ETTm2 datasets. This is because (1) the unique way IP-LLM tokenized the input time series data can yield better time series representations, and (2) the semantic space informed prompting can help further enhance the time series representation which will be further demonstrated in Section 4.5.
| Methods | IP-LLM | Time-LLM | OFA | iTransformer | Dlinear | PatchTST | TimesNet | ||||||||
| Weather | 96 | 0.050 | 0.121 | 0.147 | 0.201 | 0.162 | 0.212 | 0.253 | 0.304 | 0.176 | 0.237 | 0.149 | 0.198 | 0.172 | 0.220 |
| 192 | 0.090 | 0.164 | 0.189 | 0.234 | 0.204 | 0.248 | 0.280 | 0.319 | 0.220 | 0.282 | 0.194 | 0.241 | 0.219 | 0.261 | |
| 336 | 0.176 | 0.234 | 0.262 | 0.279 | 0.254 | 0.286 | 0.321 | 0.344 | 0.265 | 0.319 | 0.245 | 0.282 | 0.280 | 0.306 | |
| 720 | 0.300 | 0.306 | 0.304 | 0.316 | 0.326 | 0.337 | 0.364 | 0.374 | 0.333 | 0.362 | 0.314 | 0.334 | 0.365 | 0.359 | |
| Avg. | 0.154 | 0.206 | 0.225 | 0.257 | 0.237 | 0.270 | 0.304 | 0.335 | 0.248 | 0.300 | 0.225 | 0.264 | 0.259 | 0.287 | |
| Electricity | 96 | 0.136 | 0.225 | 0.131 | 0.224 | 0.139 | 0.238 | 0.147 | 0.248 | 0.140 | 0.237 | 0.129 | 0.222 | 0.168 | 0.272 |
| 192 | 0.147 | 0.240 | 0.152 | 0.241 | 0.153 | 0.251 | 0.165 | 0.267 | 0.153 | 0.249 | 0.157 | 0.240 | 0.184 | 0.289 | |
| 336 | 0.155 | 0.243 | 0.160 | 0.248 | 0.169 | 0.266 | 0.178 | 0.279 | 0.169 | 0.267 | 0.163 | 0.259 | 0.198 | 0.300 | |
| 720 | 0.183 | 0.282 | 0.192 | 0.298 | 0.206 | 0.297 | 0.322 | 0.398 | 0.203 | 0.301 | 0.197 | 0.290 | 0.220 | 0.320 | |
| Avg. | 0.155 | 0.247 | 0.158 | 0.252 | 0.167 | 0.263 | 0.203 | 0.298 | 0.166 | 0.263 | 0.161 | 0.252 | 0.192 | 0.295 | |
| Traffic | 96 | 0.347 | 0.231 | 0.362 | 0.248 | 0.388 | 0.282 | 0.367 | 0.288 | 0.410 | 0.282 | 0.360 | 0.249 | 0.593 | 0.321 |
| 192 | 0.353 | 0.240 | 0.374 | 0.247 | 0.407 | 0.290 | 0.378 | 0.293 | 0.423 | 0.287 | 0.379 | 0.256 | 0.617 | 0.336 | |
| 336 | 0.361 | 0.247 | 0.385 | 0.271 | 0.412 | 0.294 | 0.389 | 0.294 | 0.436 | 0.296 | 0.392 | 0.264 | 0.629 | 0.336 | |
| 720 | 0.392 | 0.278 | 0.430 | 0.288 | 0.450 | 0.312 | 0.401 | 0.304 | 0.466 | 0.315 | 0.432 | 0.286 | 0.640 | 0.350 | |
| Avg. | 0.363 | 0.249 | 0.388 | 0.264 | 0.414 | 0.294 | 0.389 | 0.295 | 0.433 | 0.295 | 0.390 | 0.263 | 0.620 | 0.336 | |
| ETTh1 | 96 | 0.318 | 0.377 | 0.362 | 0.392 | 0.379 | 0.402 | 0.395 | 0.420 | 0.367 | 0.396 | 0.379 | 0.407 | 0.468 | 0.475 |
| 192 | 0.389 | 0.403 | 0.398 | 0.418 | 0.415 | 0.424 | 0.427 | 0.441 | 0.401 | 0.419 | 0.428 | 0.442 | 0.484 | 0.485 | |
| 336 | 0.427 | 0.423 | 0.430 | 0.427 | 0.435 | 0.440 | 0.445 | 0.457 | 0.434 | 0.449 | 0.465 | 0.465 | 0.536 | 0.516 | |
| 720 | 0.447 | 0.458 | 0.442 | 0.457 | 0.441 | 0.459 | 0.537 | 0.530 | 0.472 | 0.493 | 0.504 | 0.500 | 0.593 | 0.537 | |
| Avg. | 0.396 | 0.416 | 0.408 | 0.423 | 0.418 | 0.431 | 0.451 | 0.462 | 0.418 | 0.439 | 0.444 | 0.453 | 0.520 | 0.505 | |
| ETTh2 | 96 | 0.245 | 0.313 | 0.268 | 0.328 | 0.289 | 0.347 | 0.304 | 0.360 | 0.301 | 0.367 | 0.296 | 0.353 | 0.376 | 0.415 |
| 192 | 0.324 | 0.365 | 0.329 | 0.375 | 0.358 | 0.392 | 0.377 | 0.403 | 0.394 | 0.427 | 0.382 | 0.404 | 0.409 | 0.440 | |
| 336 | 0.351 | 0.403 | 0.368 | 0.409 | 0.383 | 0.414 | 0.405 | 0.429 | 0.506 | 0.495 | 0.402 | 0.425 | 0.425 | 0.455 | |
| 720 | 0.363 | 0.410 | 0.372 | 0.420 | 0.438 | 0.456 | 0.443 | 0.464 | 0.805 | 0.635 | 0.444 | 0.465 | 0.488 | 0.494 | |
| Avg. | 0.320 | 0.372 | 0.334 | 0.383 | 0.367 | 0.402 | 0.382 | 0.414 | 0.502 | 0.481 | 0.381 | 0.411 | 0.425 | 0.451 | |
| ETTm1 | 96 | 0.096 | 0.211 | 0.272 | 0.334 | 0.296 | 0.353 | 0.312 | 0.366 | 0.304 | 0.348 | 0.303 | 0.351 | 0.329 | 0.377 |
| 192 | 0.186 | 0.287 | 0.310 | 0.358 | 0.335 | 0.373 | 0.347 | 0.385 | 0.336 | 0.367 | 0.341 | 0.376 | 0.371 | 0.401 | |
| 336 | 0.286 | 0.352 | 0.352 | 0.384 | 0.369 | 0.394 | 0.379 | 0.404 | 0.368 | 0.387 | 0.377 | 0.401 | 0.417 | 0.428 | |
| 720 | 0.377 | 0.411 | 0.383 | 0.411 | 0.418 | 0.424 | 0.441 | 0.442 | 0.421 | 0.418 | 0.431 | 0.436 | 0.483 | 0.464 | |
| Avg. | 0.236 | 0.315 | 0.329 | 0.372 | 0.355 | 0.386 | 0.370 | 0.399 | 0.357 | 0.389 | 0.363 | 0.391 | 0.400 | 0.417 | |
| ETTm2 | 96 | 0.082 | 0.194 | 0.161 | 0.253 | 0.170 | 0.264 | 0.179 | 0.271 | 0.168 | 0.263 | 0.173 | 0.262 | 0.201 | 0.286 |
| 192 | 0.163 | 0.263 | 0.219 | 0.293 | 0.231 | 0.306 | 0.242 | 0.313 | 0.229 | 0.310 | 0.231 | 0.300 | 0.260 | 0.329 | |
| 336 | 0.234 | 0.314 | 0.271 | 0.329 | 0.280 | 0.339 | 0.288 | 0.344 | 0.289 | 0.352 | 0.292 | 0.345 | 0.331 | 0.376 | |
| 720 | 0.337 | 0.362 | 0.352 | 0.379 | 0.373 | 0.402 | 0.378 | 0.397 | 0.416 | 0.437 | 0.371 | 0.394 | 0.428 | 0.430 | |
| Avg. | 0.204 | 0.283 | 0.251 | 0.313 | 0.265 | 0.328 | 0.272 | 0.331 | 0.275 | 0.340 | 0.267 | 0.325 | 0.305 | 0.355 | |
@@ 4.2 Short-term Forecasting
Setup. We also evaluate the effectiveness of IP-LLM with the short-term forecasting setting based on the M4 datasets (Makridakis et al., 2018). It contains a collection of marketing data that are sampled at different frequencies. Details of the datasets can be found in Appendix A.3. The prediction horizons are significantly shorter than the long-term forecasting setting and are set as [6,48]. The input lengths are twice the prediction horizons, similar to the experiment setting in (Jin et al., 2024; Tian Zhou & Jin, 2023). The evaluation metrics for short-term forecasting are symmetric mean absolute percentage error (SMAPE), mean absolute scaled error (MSAE), and overall weighted average (OWA). The details of these evaluation metrics are provided in Appendix A.4.
Results. Table 2 summarizes the short-term forecasting results and the full experiment results are shown in Appendix Appendix C, Table 9. We observe that IP-LLM outperforms all other baselines by a large margin and is slightly better than Time-LLM. In terms of OWA, IP-LLM improves Time-LLM and OFA with 3.5% and 13.2 %, respectively. This could attribute to the tokenization design as well as the semantic space informed prompting within IP-LLM.
| Methods | IP-LLM | Time-LLM | OFA | iTransformer | Dlinear | PatchTST | TimesNet | FEDformer | Autoformer | |
| Avg. | SMAPE | 11.832 | 11.983 | 12.690 | 12.142 | 13.639 | 12.059 | 12.880 | 13.160 | 12.909 |
| MASE | 1.584 | 1.595 | 1.808 | 1.631 | 2.095 | 1.623 | 1.836 | 1.775 | 1.771 | |
| OWA | 0.830 | 0.859 | 0.94 | 0.874 | 1.051 | 0.869 | 0.955 | 0.949 | 0.939 | |
@@ 4.3 Few-shot Forecasting
| Methods | IP-LLM | Time-LLM | OFA | iTransformer | Dlinear | PatchTST | TimesNet | |||||||
| Weather | 0.228 | 0.268 | 0.234 | 0.273 | 0.238 | 0.275 | 0.308 | 0.338 | 0.241 | 0.283 | 0.242 | 0.279 | 0.279 | 0.301 |
| Electricity | 0.178 | 0.273 | 0.175 | 0.270 | 0.176 | 0.269 | 0.196 | 0.293 | 0.180 | 0.280 | 0.180 | 0.273 | 0.323 | 0.392 |
| Traffic | 0.426 | 0.305 | 0.429 | 0.306 | 0.440 | 0.310 | 0.495 | 0.361 | 0.447 | 0.313 | 0.430 | 0.305 | 0.951 | 0.535 |
| ETTh1 | 0.565 | 0.524 | 0.556 | 0.522 | 0.590 | 0.525 | 0.910 | 0.860 | 0.691 | 0.600 | 0.633 | 0.542 | 0.869 | 0.628 |
| ETTh2 | 0.366 | 0.392 | 0.370 | 0.394 | 0.397 | 0.421 | 0.489 | 0.483 | 0.605 | 0.538 | 0.415 | 0.431 | 0.479 | 0.465 |
| ETTm1 | 0.402 | 0.422 | 0.404 | 0.427 | 0.464 | 0.441 | 0.728 | 0.565 | 0.411 | 0.429 | 0.501 | 0.466 | 0.677 | 0.537 |
| ETTm2 | 0.140 | 0.242 | 0.177 | 0.261 | 0.188 | 0.269 | 0.245 | 0.322 | 0.213 | 0.303 | 0.191 | 0.274 | 0.212 | 0.285 |
| Methods | IP-LLM | Time-LLM | OFA | iTransformer | Dlinear | PatchTST | TimesNet | |||||||
| Weather | 0.257 | 0.295 | 0.260 | 0.309 | 0.263 | 0.301 | 0.309 | 0.339 | 0.263 | 0.308 | 0.269 | 0.303 | 0.298 | 0.318 |
| Electricity | 0.186 | 0.281 | 0.179 | 0.268 | 0.178 | 0.273 | 0.201 | 0.296 | 0.176 | 0.275 | 0.181 | 0.277 | 0.402 | 0.453 |
| Traffic | 0.419 | 0.298 | 0.423 | 0.298 | 0.434 | 0.305 | 0.450 | 0.324 | 0.450 | 0.317 | 0.418 | 0.296 | 0.867 | 0.493 |
| ETTh1 | 0.642 | 0.546 | 0.627 | 0.543 | 0.681 | 0.560 | 1.070 | 0.710 | 0.750 | 0.611 | 0.694 | 0.569 | 0.925 | 0.647 |
| ETTh2 | 0.380 | 0.415 | 0.382 | 0.418 | 0.400 | 0.433 | 0.488 | 0.475 | 0.694 | 0.577 | 0.827 | 0.615 | 0.439 | 0.448 |
| ETTm1 | 0.416 | 0.421 | 0.425 | 0.434 | 0.472 | 0.450 | 0.784 | 0.596 | 0.400 | 0.417 | 0.526 | 0.476 | 0.717 | 0.561 |
| ETTm2 | 0.279 | 0.325 | 0.274 | 0.323 | 0.308 | 0.346 | 0.356 | 0.388 | 0.399 | 0.426 | 0.314 | 0.352 | 0.344 | 0.372 |
@@ 4.4 Ablation Studies and Parameter Sensitivity
Ablation Studies. We conduct Ablation Studies on the ETTh1 and ETTm1 datasets to validate the effectiveness of different components for IP-LLM. Table 5.A reports the forecasting results when decomposition is deactivated and substituted exclusively with patching, as well as the results when the alignment and prompting features are disabled. Without the decomposition component, the performance of IP-LLM significantly deteriorates. This suggests that the decomposition contributes substantially to enrich the time series representation within IP-LLM. We also observe the performance of IP-LLM drops when we deactivate the prompting and alignment component, which also implies the importance of this module IP-LLM.
| Ablation Setting | Long-term Forecasting | |||
| ETTh1-96 | ETTh1-192 | ETTm1-96 | ETTm1-192 | |
| Default | 0.318 | 0.389 | 0.096 | 0.186 |
| A.1 w/o Decomposition | 0.396 | 0.457 | 0.203 | 0.298 |
| A.2 w/o Prompt & Alignment | 0.346 | 0.395 | 0.113 | 0.204 |
| B.1 Prompt Length (k) = 2 | 0.327 | 0.404 | 0.100 | 0.190 |
| B.2 Prompt Length (k) = 4 | 0.336 | 0.416 | 0.103 | 0.193 |
| B.3 Prompt Length (k) = 8 (default) | 0.318 | 0.389 | 0.096 | 0.186 |
| B.4 Prompt Length (k) = 16 | 0.336 | 0.409 | 0.108 | 0.195 |
| B.5 Prompt Length (k) = 32 | 0.338 | 0.427 | 0.114 | 0.205 |
| C.1 = 0.0 | 0.323 | 0.407 | 0.107 | 0.194 |
| C.2 = 0.01 | 0.320 | 0.391 | 0.101 | 0.191 |
| C.3 = 0.05 (default) | 0.318 | 0.389 | 0.096 | 0.186 |
| C.4 = 0.1 | 0.321 | 0.393 | 0.103 | 0.192 |
| C.5 = 1 | 0.325 | 0.408 | 0.110 | 0.198 |
| D.1 = 50 | 0.325 | 0.410 | 0.110 | 0.204 |
| D.2 = 100 | 0.324 | 0.408 | 0.108 | 0.203 |
| D.3 = 500 | 0.320 | 0.393 | 0.101 | 0.193 |
| D.4 = 1000 (default) | 0.318 | 0.389 | 0.096 | 0.186 |
| D.5 = 2000 | 0.315 | 0.385 | 0.091 | 0.183 |
| D.6 = 5000 | 0.314 | 0.383 | 0.089 | 0.180 |
Parameter Sensitivity. Table 5.B presents the experiment results when the length of the prompt varies. Within a limited range, i.e. 2 to 8, an increase in the prompt length tends to improve the forecasting performance. However, excessive prompt length, such as lengths of 16 or 32, results in a significant decline in the forecasting accuracy. A similar pattern can be observed in the hyperparameter analysis of , which controls the strength of alignment. As shown in Table 5.C, when varies from 0 to 0.05, slightly larger is beneficial for representation learning within the joint space, showing better forecasting results. On the other hand, larger tends to lead to indistinguishable time series representation and the forecasting performance will decrease. Finally, Table 5.D shows the effects of choosing different numbers of semantic anchors. Generally, an increased number of semantic anchors improves the forecasting results. We maintain the size of as 1000 in our experiments to ensure computational efficiency. We conjecture that the small number hinders the learning of highly representative semantic anchors in the joint space and thus will generate less informed prompts for time series embedding. We visualize the prompted time series embeddings with the different number of semantic anchors in Appendix E, Figure 5. We notice that a smaller quantity of semantic anchors leads to a less dispersed distribution in the joint space, indicating that the generated prompts could be less informative for time series embedding.

### v2

@@ 4.1 Long-term Forecasting
Setup. For long-term forecasting, we evaluate the effectiveness of IP-LLM on Weather, Electricity, Traffic, and four ETT datasets (i.e., ETTh1, ETTh2, ETTm1, and ETTm2), which have been widely adopted as benchmarking datasets for long-term forecasting tasks. Details of these datasets are shown in Appendix A.3, Table 5. The input time series length is 512, and we evaluate the performance on four different horizons . The evaluation metrics include the mean square error (MSE) and the mean absolute error (MAE).
@@
Results. We compare the forecasting results of IP-LLM to 6 selected baselines in Table 1. Due to the space limitation, the comparisons with the other 6 baselines are provided in Appendix B and Table 6. We can observe that LLMs based forecasting methods, i.e., Time-LLM and OFA, generally achieve better performance than other baseline methods. This should be attributed to the prevalent expressibility of LLMs and their associated prompt-tuning and fine-tuning strategies, respectively. Moreover, most of the time, IP-LLM outperforms Time-LLM and OFA over 7 different datasets. This is because (1) the unique way IP-LLM tokenized the input time series data can yield better time series representations, and (2) the semantic space informed prompting can help further enhance the time series representation which will be further demonstrated in Section 4.5.
@@ 4.2 Short-term Forecasting
Setup. We also evaluate the effectiveness of IP-LLM with the short-term forecasting setting based on the M4 datasets (Makridakis et al., 2018). It contains a collection of marketing data that are sampled at different frequencies. Details of the datasets can be found in Appendix A.3. The prediction horizons are significantly shorter than the long-term forecasting setting and are set between 6 and 48. The input lengths are twice the prediction horizons, similar to the experiment setting in (Jin et al., 2024; Tian Zhou & Jin, 2023). The evaluation metrics for short-term forecasting are symmetric mean absolute percentage error (SMAPE), mean absolute scaled error (MASE), and overall weighted average (OWA). The details of these evaluation metrics are provided in Appendix A.4.
Results. Table 2 summarizes the short-term forecasting results and the full experiment results are shown in Appendix Appendix C, Table 7. We observe that IP-LLM outperforms all other baselines by a large margin and is slightly better than PatchTST. This could attribute to the tokenization design as well as the semantic space informed prompting within IP-LLM.
| Methods | IP-LLM | Time-LLM(G) | OFA | iTransformer | Dlinear | PatchTST | TimesNet | FEDformer | Autoformer | |
| Avg. | SMAPE | 12.021 | 12.494 | 12.690 | 12.142 | 13.639 | 12.059 | 12.880 | 13.160 | 12.909 |
| MASE | 1.612 | 1.731 | 1.808 | 1.631 | 2.095 | 1.623 | 1.836 | 1.775 | 1.771 | |
| OWA | 0.857 | 0.913 | 0.94 | 0.874 | 1.051 | 0.869 | 0.955 | 0.949 | 0.939 | |
| Methods | IP-LLM | Time-LLM(G) | OFA | iTransformer | Dlinear | PatchTST | TimesNet | |||||||
| Weather | 0.233 | 0.272 | 0.237 | 0.275 | 0.238 | 0.275 | 0.308 | 0.338 | 0.241 | 0.283 | 0.242 | 0.279 | 0.279 | 0.301 |
| Electricity | 0.175 | 0.271 | 0.177 | 0.273 | 0.176 | 0.269 | 0.196 | 0.293 | 0.180 | 0.280 | 0.180 | 0.273 | 0.323 | 0.392 |
| Traffic | 0.427 | 0.307 | 0.429 | 0.307 | 0.440 | 0.310 | 0.495 | 0.361 | 0.447 | 0.313 | 0.430 | 0.305 | 0.951 | 0.535 |
| ETTh1 | 0.593 | 0.529 | 0.785 | 0.553 | 0.590 | 0.525 | 0.910 | 0.860 | 0.691 | 0.600 | 0.633 | 0.542 | 0.869 | 0.628 |
| ETTh2 | 0.419 | 0.439 | 0.424 | 0.441 | 0.397 | 0.421 | 0.489 | 0.483 | 0.605 | 0.538 | 0.415 | 0.431 | 0.479 | 0.465 |
| ETTm1 | 0.455 | 0.435 | 0.487 | 0.461 | 0.464 | 0.441 | 0.728 | 0.565 | 0.411 | 0.429 | 0.501 | 0.466 | 0.677 | 0.537 |
| ETTm2 | 0.284 | 0.332 | 0.305 | 0.344 | 0.293 | 0.335 | 0.336 | 0.373 | 0.316 | 0.368 | 0.296 | 0.343 | 0.320 | 0.353 |
| Methods | IP-LLM | Time-LLM(G) | OFA | iTransformer | Dlinear | PatchTST | TimesNet | |||||||
| Weather | 0.260 | 0.297 | 0.264 | 0.301 | 0.263 | 0.301 | 0.309 | 0.339 | 0.263 | 0.308 | 0.269 | 0.303 | 0.298 | 0.318 |
| Electricity | 0.179 | 0.275 | 0.181 | 0.279 | 0.178 | 0.273 | 0.201 | 0.296 | 0.176 | 0.275 | 0.181 | 0.277 | 0.402 | 0.453 |
| Traffic | 0.420 | 0.299 | 0.423 | 0.302 | 0.434 | 0.305 | 0.450 | 0.324 | 0.450 | 0.317 | 0.418 | 0.296 | 0.867 | 0.493 |
| ETTh1 | 0.650 | 0.550 | 0.891 | 0.627 | 0.681 | 0.560 | 1.070 | 0.710 | 0.750 | 0.611 | 0.694 | 0.569 | 0.925 | 0.647 |
| ETTh2 | 0.380 | 0.413 | 0.581 | 0.519 | 0.400 | 0.433 | 0.488 | 0.475 | 0.694 | 0.577 | 0.827 | 0.615 | 0.439 | 0.448 |
| ETTm1 | 0.455 | 0.446 | 0.524 | 0.479 | 0.472 | 0.450 | 0.784 | 0.596 | 0.400 | 0.417 | 0.526 | 0.476 | 0.717 | 0.561 |
| ETTm2 | 0.296 | 0.342 | 0.325 | 0.361 | 0.308 | 0.346 | 0.356 | 0.388 | 0.399 | 0.426 | 0.314 | 0.352 | 0.344 | 0.372 |
@@ 4.4 Ablation Studies and Parameter Sensitivity
We conduct ablation studies on the ETTh2 and ETTm2 datasets to evaluate the parameter sensitivity for IP-LLM. Figure 3 (1) and (4) presents the experiment results when the length of the prompt varies on ETTh2 and ETTm2, respectively. Within a limited range, i.e. 2 to 8, an increase in the prompt length tends to improve the forecasting performance. However, excessive prompt length, such as lengths of 16 or 32, results in a significant decline in the forecasting accuracy. A similar pattern can be observed in the hyperparameter analysis of the , which controls the strength of alignment. As shown in Figure 3 (2) and (5), when varies from 0 to 0.05, slightly larger is beneficial for representation learning within the joint space, showing better forecasting results. On the other hand, larger tends to lead to indistinguishable time series representation and the forecasting performance will thus decrease. Finally, Figure 3 (3) and (6) show the effects of choosing different numbers of semantic anchors. Generally, an increased number of semantic anchors improves the forecasting results. We conjecture that the small number hinders the learning of highly representative semantic anchors in the joint space and thus will generate less informed prompts for time series embedding. We visualize the prompted time series embeddings with different numbers of semantic anchors in Appendix E, Figure 6. We notice that a smaller quantity of semantic anchors leads to a less dispersed distribution in the joint space, indicating that the generated prompts could be less informative for time series embedding. We also perform ablation studies by incrementally adding the “alignment & prompting” and “decomposition” modules. In Appendix E Table 10, we observe the forecasting performance increases when we sequentially activate the prompting & alignment component and the decomposition component, which implies the importance of these modules in IP-LLM.